The sum of two numbers is 94. The second is 6 more than 3 times the first. What are the two numbers?
Answer by elima(1433)

The sum of two numbers is 94. The second is 6 more than 3 times the first. What are the two numbers?
x+y=94
y+6=3x
------------
x+(3x-6)=94
4x=100
x=25
-------------
25+y=94
y=94-25
y=69
:)